If angle 1 and 2 are supplementary angles and angle 1 is 60 degrees less than angle 2, what is the measurment of each angle?
Answer by rfer(16322)

x+x-60=180
2x=180+60
x=240/2
x=120 (angle 2)
x-60=60 (angle 1)